What is the [H2CO3], [HCO3-] and pH of a buffer solution which is composed of 25 mL of 0.05 M H2CO3 and 475 mL of 0.263 M NaHCO3? (Ka for H2CO3 = 4.2 x 10-7).


1
Expert's answer
2020-07-13T07:23:08-0400

Solution:

V(buffer solution) = V(H2CO3) + V(NaHCO3) = 25 mL + 475 mL = 500 mL


[H2CO3] = C(H2CO3) × V(H2CO3) / V(solution) = (0.05 M) × (25 mL / 500 mL) = 0.0025 M

[H2CO3] = 0.0025 M


[HCO3-] = C(NaHCO3) × V(NaHCO3) / V(solution) = (0.263 M) × (475 mL / 500 mL) = 0.24985 M = 0.25 M

[HCO3-] = 0.2500 M


To calculate pH of a buffer solution Henderson-Hasselbalch Equation can be used.

Henderson-Hasselbalch Equation can be expressed as:

[H+] = Ka × ([HA]/[A-])

take –log of both sides of equation

pH = pKa + log([A-]/[HA])

[A-] = [HCO3-]

[HA] = [H2CO3]

pH = pKa + log([HCO3-] / [H2CO3])


pKa = -log(Ka) = -log(4.2×10-7) = 6.376 = 6.38

Thus:

pH = 6.38 + log(0.2500 / 0.0025) = 6.38 + 2 = 8.38

pH = 8.38


Answer:

[H2CO3] = 0.0025 M

[HCO3-] = 0.2500 M

pH = 8.38